Hello! Can you please help me answer this question right away?
A confectioner finds that it is hard to sell chocolate worth P28.80 a lb which he has 6 lbs. He mixes them with a candy he usually sells at P16.80 a lb and offers the resulting mixture at P21.60 a lb. How much of P16.80 of candy should he use?
*P - pesos
*lb - pounds
Thank You.

Answer by mananth(16949)

------------------------ price ---------------- quantity
Chocolate type I 28.8 ---------------- 6 pounds
Candy 16.8 ---------------- x pounds
Mixture 21.6 ---------------- 6 + x pounds
The sum of idividual costs of the two types = cost of mixture
28.8 * 6 + 16.8 x = 21.6 (6+x)
172.8 + 16.8 x = 129.6 + 21.6 x
16.8 x -21.6 x= 129.6 -172.8
-4.8 x = -43.2
/ -4.8
x= 9 pounds of Candy

m.ananth@hotmail.ca